OpenStudy (anonymous):

can someone help me find an equation for the inverse of the function. f(x)=2x - 3^x/4

OpenStudy (irishboy123):

lmLtfy :p \(\large f(x)=2x - \frac{3^x}{4}\) ??

OpenStudy (anonymous):

Yes that's it

OpenStudy (irishboy123):

i don't think you can do it because as soon as you try to get that x out of the exponential, you lock the other x in a log do you know if there is an answer and what that answer is? if so i can try to hack it or you can bump and wait for someone else to come along to complicate matters, the other thing is this fails the so-called Horizontal Line Test - there is a local maximum, as you can see from the attached, so for f(x), 2 x's can map to the same y; but the inverse would require the opposite, ie that 1 y maps to 2 x's so there is no function so that you would need to restrict the domain to the LHS or RHS of the turning point .... not that helpful, i know :p

OpenStudy (anonymous):

Wel the the multiple choice answers provided for this question are: A. f^-1(x)=3x-2 B. f^-1(x)=- 3/4x+2 C. f^-11(x)=4/5x D. f^-1(x)=4/5x+3/4 I'm just not sure of the answer.

OpenStudy (irishboy123):

none of these give you an inverse to what you posted maybe you should post the actual question or a screenshot. i think that could help greatly.

OpenStudy (irishboy123):

that is think is "your" gmail account ! just a picture, a screen grab , might be really informative, thx.